Balance the following chemical equations. Always show your solution.
1. ____ N 2 O 5  ____ N 2 + ____O 2

2. ____ C 5 H 12 + ____ O 2  ____ CO 2 + ____ H 2 O

3. ____ C 4 H 6 O 3 + ____ H 2 O  ____ C 2 H 4 O 2
1
Expert's answer
2020-05-03T15:09:50-0400

Solution:

(1)

First, write the unbalanced equation: N2O5 = N2 + O2 (unbalanced)

Next, count the number of each type of atom present in the unbalanced equation.



Though nitrogen is balanced, changes in coefficients are needed to balance the number of oxygen atoms. To balance the number of oxygen atoms, a reasonable first attempt would be to change the coefficients for the O2 and N2O5 to integers that will yield 10 O atoms (the least common multiple for the O atom subscripts in these two formulas).

2N2O5 = N2 + 5O2 (unbalanced)



The N atom balance has been upset by this change; it is restored by changing the coefficient for the reactant N2 to 2.

2N2O5 = 2N2 + 5O2


The numbers of N and O atoms on either side of the equation are now equal, and so the equation is balanced.

The balanced chemical equation (1): 2N2O5 = 2N2 + 5O2



(2)

First, write the unbalanced equation: C5H12 + O2 = CO2 + H2O (unbalanced)

Next, count the number of each type of atom present in the unbalanced equation.



To balance the number of hydrogen atoms, a reasonable first attempt would be to change the coefficients for the C5H12 and H2O to integers that will yield 12 H atoms (the least common multiple for the H atom subscripts in these two formulas).

C5H12 + O2 = CO2 + 6H2O (unbalanced)



To balance the number of carbon atoms, a reasonable attempt would be to change the coefficients for the C5H12 and CO2 to integers that will yield 5 C atoms (the least common multiple for the C atom subscripts in these two formulas).

C5H12 + O2 = 5CO2 + 6H2O (unbalanced)



The O atom balance is restored by changing the coefficient for the reactant O2 to 8.

C5H12 + 8O2 = 5CO2 + 6H2O



The numbers of C, H and O atoms on either side of the equation are now equal, and so the equation is balanced.

The balanced chemical equation (2): C5H12 + 8O2 = 5CO2 + 6H2O



(3)

First, write the unbalanced equation: C4H6O3 + H2O = C2H4O2 (unbalanced)

Next, count the number of each type of atom present in the unbalanced equation.


To balance the number of carbon atoms, a reasonable first attempt would be to change the coefficients for the C4H6O3 and C2H4O2 to integers that will yield 4 C atoms (the least common multiple for the C atom subscripts in these two formulas).

C4H6O3 + H2O = 2C2H4O2


The numbers of C, H and O atoms on either side of the equation are now equal, and so the equation is balanced.

The balanced chemical equation (3): C4H6O3 + H2O = 2C2H4O2